# The Angular Momentum Penrose Inequality

### v2

@@ 5.1 The Conformal Equation
We give a detailed proof using Lockhart–McOwen theory, with careful treatment of the marginally stable case.
Step 1: Asymptotic structure of the operator. By Theorem 4.12(iii), the Jang metric $\bar{g}$ converges exponentially to the cylindrical metric $dt^{2}+g_{\Sigma}$ on the ends:
$\bar{g}=dt^{2}+g_{\Sigma}+O(e^{-\beta_{0}t})\quad\text{as }t\to\infty,$
where the exponential decay rate $\beta_{0}>0$ is determined by the spectral gap of the MOTS stability operator. Specifically:
  • For strictly stable MOTS ($\lambda_{1}(L_{\Sigma})>0$): $\beta_{0}=2\sqrt{\lambda_{1}(L_{\Sigma})}$, where $\lambda_{1}$ is the principal eigenvalue of the stability operator (14).
  • For marginally stable MOTS ($\lambda_{1}(L_{\Sigma})=0$): $\beta_{0}=2$, arising from the subleading spectral term. This is the borderline case discussed in Step 4 below.
This relationship between $\beta_{0}$ and stability follows from the eigenvalue problem for the linearized Jang operator at the MOTS; see [8, Proposition 3.4] and [29, Section 4].
The Lockhart–McOwen theory [36] applies to operators of the form $L=L_{\infty}+Q$ where:
  • $L_{\infty}=-8(\partial_{t}^{2}+\Delta_{\Sigma})+R_{\Sigma}$ is the translation-invariant limit operator on the exact cylinder $\mathbb{R}\times\Sigma$.
  • $Q$ is a perturbation satisfying $|Q|=O(e^{-\beta_{0}t})$ as $t\to\infty$, arising from the deviation $\bar{g}-(dt^{2}+g_{\Sigma})$.
Step 2: Indicial roots computation. The indicial roots are determined by seeking solutions of $L_{\infty}\psi=0$ of the form $\psi(t,y)=e^{\gamma t}\varphi(y)$ where $\varphi$ is an eigenfunction of the cross-sectional operator. Substituting:
$L_{\infty}(e^{\gamma t}\varphi)=e^{\gamma t}\bigl(-8\gamma^{2}\varphi-8\Delta_{\Sigma}\varphi+R_{\Sigma}\varphi\bigr)=0.$
Thus $\varphi$ must be an eigenfunction of $8\Delta_{\Sigma}+R_{\Sigma}$ on $(\Sigma,g_{\Sigma})$:
$(-8\Delta_{\Sigma}+R_{\Sigma})\varphi=\lambda\varphi,$
and the indicial root satisfies $8\gamma^{2}+\lambda=0$, giving:
$\gamma=\pm\sqrt{\lambda/8}.$
Step 3: Eigenvalue lower bound for the cross-sectional operator. We need to show that the operator $8\Delta_{\Sigma}+R_{\Sigma}$ on $(\Sigma,g_{\Sigma})\cong S^{2}$ has strictly positive principal eigenvalue $\lambda_{0}>0$.
Claim: For any Riemannian metric on $S^{2}$ with scalar curvature $R_{\Sigma}$ satisfying $\int_{\Sigma}R_{\Sigma}\,d\sigma=8\pi$ (by Gauss–Bonnet), the operator $8\Delta_{\Sigma}+R_{\Sigma}$ has $\lambda_{0}>0$.
Proof of Claim.
The principal eigenvalue is given by the variational formula:
$\lambda_{0}=\inf_{\|\varphi\|_{L^{2}}=1}\int_{\Sigma}\bigl(8|\nabla\varphi|^{2}+R_{\Sigma}\varphi^{2}\bigr)d\sigma.$
We show $\lambda_{0}>0$ by contradiction. Suppose $\lambda_{0}\leq 0$. Then there exists $\varphi\in W^{1,2}(\Sigma)$ with $\|\varphi\|_{L^{2}}=1$ and
$\int_{\Sigma}\bigl(8|\nabla\varphi|^{2}+R_{\Sigma}\varphi^{2}\bigr)d\sigma\leq 0.$
Case 1: $\lambda_{0}=0$ with eigenfunction $\varphi_{0}$. If $\lambda_{0}=0$ is achieved by an eigenfunction $\varphi_{0}$, then:
$8\Delta_{\Sigma}\varphi_{0}+R_{\Sigma}\varphi_{0}=0.$
Integrating over $\Sigma$:
$\int_{\Sigma}R_{\Sigma}\varphi_{0}\,d\sigma=8\int_{\Sigma}\Delta_{\Sigma}\varphi_{0}\,d\sigma=0$
by the divergence theorem on a closed surface. Since $\varphi_{0}$ is an eigenfunction with $\lambda_{0}=0$, it cannot change sign (principal eigenfunctions are either strictly positive or strictly negative). WLOG, assume $\varphi_{0}>0$. Then:
$\int_{\Sigma}R_{\Sigma}\varphi_{0}\,d\sigma=0\quad\text{with }\varphi_{0}>0.$
This implies $R_{\Sigma}$ must change sign on $\Sigma$ (otherwise the integral would be strictly positive or negative).
Now, use the constraint from stability. For a stable MOTS $\Sigma$, the Galloway–Schoen theorem [27] implies that $\Sigma$ has non-negative Gaussian curvature: $K_{\Sigma}\geq 0$ everywhere. Since $R_{\Sigma}=2K_{\Sigma}$ for surfaces, this means $R_{\Sigma}\geq 0$ on $\Sigma$. Combined with $\int_{\Sigma}R_{\Sigma}=8\pi>0$, we have $R_{\Sigma}\geq 0$ (not identically zero). Therefore:
$\int_{\Sigma}R_{\Sigma}\varphi_{0}\,d\sigma>0\quad\text{(since }R_{\Sigma}\geq 0,\varphi_{0}>0,\text{ and }R_{\Sigma}\not\equiv 0),$
contradicting $\int_{\Sigma}R_{\Sigma}\varphi_{0}=0$.
Case 2: $\lambda_{0}<0$. This would require $\int_{\Sigma}R_{\Sigma}\varphi^{2}<0$ for some $\varphi$ (since $\int 8|\nabla\varphi|^{2}\geq 0$). But $R_{\Sigma}\geq 0$ on a stable MOTS, so $\int R_{\Sigma}\varphi^{2}\geq 0$ for all $\varphi$, contradiction.
Conclusion: $\lambda_{0}>0$ for the operator $8\Delta_{\Sigma}+R_{\Sigma}$ on a stable MOTS. ∎
Remark: The key input is that stable MOTS in data satisfying DEC have $R_{\Sigma}=2K_{\Sigma}\geq 0$ (non-negative Gaussian curvature). This follows from the stability inequality and the Gauss–Bonnet–based argument in [27]. For unstable MOTS, $R_{\Sigma}$ can be negative somewhere, and the argument fails.
Since $\lambda_{0}>0$, the smallest indicial root satisfies $|\gamma_{0}|=\sqrt{\lambda_{0}/8}>0$.
Step 4: Treatment of the marginally stable case. The MOTS stability operator $\mathcal{L}_{\Sigma}$ may have $\lambda_{1}=0$ (marginal stability), but this is distinct from the operator $8\Delta_{\Sigma}+R_{\Sigma}$ appearing in the Lichnerowicz equation. The key observation is:
  1. 1.
    The MOTS stability operator $\mathcal{L}_{\Sigma}$ governs deformations of $\Sigma$ as a trapped surface.
  2. 2.
    The Lichnerowicz operator $8\Delta_{\Sigma}+R_{\Sigma}$ governs the conformal factor on the cylindrical end.
  3. 3.
    These are different operators; marginal MOTS stability ($\lambda_{1}(\mathcal{L}_{\Sigma})=0$) does not imply $\lambda_{0}(-8\Delta_{\Sigma}+R_{\Sigma})=0$.
In fact, for any stable MOTS $\Sigma\cong S^{2}$, we have shown $\lambda_{0}(-8\Delta_{\Sigma}+R_{\Sigma})>0$ regardless of whether the MOTS stability eigenvalue is zero or positive.
Step 5: Fredholm index computation. By [36, Theorem 1.1], the operator $L:W^{2,2}_{\beta}\to L^{2}_{\beta}$ is Fredholm if and only if $\beta$ is not an indicial root. The Fredholm index is:
$\text{ind}(L)=-\sum_{\gamma:\,0<\gamma<\beta}m(\gamma)+\sum_{\gamma:\,\beta<\gamma<0}m(\gamma),$
where $m(\gamma)$ is the multiplicity of the indicial root $\gamma$.
Since the smallest positive indicial root satisfies $\gamma_{0}=\sqrt{\lambda_{0}/8}>0$, for $\beta\in(-\gamma_{0},0)$:
  • The interval $(0,\beta)$ contains no indicial roots (since $\beta<0$).
  • The interval $(\beta,0)$ contains no indicial roots (since $\gamma_{0}>0>\beta$).
Therefore $\text{ind}(L)=0$.
Step 6: Injectivity. To show $L$ is an isomorphism (not just Fredholm of index zero), we verify $\ker(L)=\{0\}$ on $W^{2,2}_{\beta}$.
Suppose $Lv=0$ with $v\in W^{2,2}_{\beta}$. Since $\beta<0$, we have $v\to 0$ as $t\to\infty$ (along the cylindrical end). Multiplying by $v$ and integrating:
$\int_{\bar{M}}\bigl(8|\nabla v|^{2}+R_{\bar{g}}v^{2}\bigr)dV_{\bar{g}}=0.$
By the Bray–Khuri identity, $R_{\bar{g}}\geq 0$ on the Jang manifold (under DEC). Therefore each term is non-negative, forcing $\nabla v=0$ and (where $R_{\bar{g}}>0$) $v=0$. Combined with the boundary condition $v\to 0$, the maximum principle implies $v\equiv 0$.
Therefore $L$ is injective, and being Fredholm of index zero, it is an isomorphism. ∎
@@
Let $(\bar{M},\bar{g})$ be the Jang manifold from Theorem 4.12 with cylindrical end $\mathcal{C}\cong[0,\infty)\times\Sigma$. Let $R_{\bar{g}}\geq 0$ be the scalar curvature (guaranteed by DEC via the Bray–Khuri identity) and $\Lambda_{J}=\frac{1}{8}|\sigma^{TT}|^{2}_{\bar{g}}\geq 0$ the TT-contribution. Then the AM-Lichnerowicz equation (40) admits a unique solution $\phi\in C^{2,\alpha}(\bar{M})\cap C^{0}(\overline{\bar{M}})$ satisfying:
@@
    Horizon normalization: $\phi|_{\Sigma}=1$, interpreted as $\lim_{t\to\infty}\phi(t,y)=1$ along the cylindrical end;
@@
    Asymptotic normalization: $\phi(x)\to 1$ as $|x|\to\infty$ in the asymptotically flat end, with decay $|\phi-1|=O(r^{-\tau})$;
@@
    Strict positivity: $\phi>0$ throughout $\bar{M}$, with $\inf_{\bar{M}}\phi>0$;
  4. (iv)
    Exponential convergence on cylinder: On the cylindrical end, $|\phi(t,y)-1|\leq Ce^{-\kappa t}$ where $\kappa=\min(\gamma_{0},\beta_{0})>0$ with $\gamma_{0}=\sqrt{\lambda_{0}/8}$ from Lemma 5.5 and $\beta_{0}$ from Theorem 4.12(iii).
Key Consequence: The conformal metric $\tilde{g}=\phi^{4}\bar{g}$ has non-negative scalar curvature:
$R_{\tilde{g}}=\Lambda_{J}\phi^{-12}\geq 0,$
with strict positivity where the data has non-trivial rotational contribution ($\Lambda_{J}>0$). This is the crucial property enabling the AMO monotonicity argument.
Remark 5.7 (Critical Clarification: Why the Supersolution Issue is Not an Obstacle).
A potential concern is whether the conformal factor $\phi$ satisfies $\phi\leq 1$, which would follow from the naive supersolution argument if $R_{\bar{g}}\geq\Lambda_{J}$. We provide a complete resolution showing this bound is not required for the main result:
  1. 1.
    The monotonicity formula (Theorem 6.22) requires only $R_{\tilde{g}}\geq 0$. By the conformal transformation formula and the AM-Lichnerowicz equation, $R_{\tilde{g}}=\Lambda_{J}\phi^{-12}\geq 0$ holds automatically for any positive solution $\phi>0$, regardless of whether $\phi\leq 1$ or $\phi>1$.
  2. 2.
    The mass inequality $M_{\mathrm{ADM}}(\tilde{g})\leq M_{\mathrm{ADM}}(g)$ follows from a direct energy argument. We establish this in Lemma 5.11 using the structure of the Jang–conformal construction, without requiring $\phi\leq 1$.
  3. 3.
[…]
Summary of the logical chain (independent of $\phi\leq 1$):
@@
    DEC on $(M,g,K)$ $\Rightarrow$ $R_{\bar{g}}\geq 0$ on Jang manifold (Bray–Khuri identity);

### v3

@@ 5.1 The Conformal Equation
We verify the key claims.
Step 1: Intrinsic definition of $(E,B)$. The formulas for $E_{ij}$ and $B_{ij}$ involve only:
  • The Ricci tensor $R_{ij}$ (determined by $g$);
  • The extrinsic curvature $K_{ij}$ (given data);
  • The Levi-Civita connection $\nabla$ of $g$.
No embedding into a spacetime is required—these are the Gauss–Codazzi projections of the spacetime Weyl tensor onto the initial surface, but computed intrinsically. For $(g,K)\in C^{k,\beta}_{-\tau}\times C^{k-1,\beta}_{-\tau-1}$ with $k\geq 3$, we have $(E,B)\in C^{k-2,\beta}_{-\tau-2}(M;S^{2}_{0}T^{*}M)$.
Step 2: Asymptotic matching. For asymptotically flat data with ADM mass $M$ and Komar angular momentum $J$, the reference Kerr tensors are:
$E^{\mathrm{Kerr}}_{ij}$ $=\frac{M}{r^{3}}\left(\delta_{ij}-3\hat{r}_{i}\hat{r}_{j}\right)+\frac{3Ma^{2}}{r^{5}}\left(\text{spin-2 harmonics}\right)+O(r^{-6}),$
$B^{\mathrm{Kerr}}_{ij}$ $=\frac{3Ma}{r^{4}}\epsilon_{(i}{}^{kl}\hat{r}_{j)}\hat{r}_{k}\hat{z}_{l}+O(r^{-5}),$
where $a=J/M$, $\hat{r}=x/r$, and $\hat{z}$ is the symmetry axis unit vector. The Kerr deviation tensor is then:
$\mathcal{S}_{(g,K),ij}=(E_{ij}-E^{\mathrm{Kerr}}_{ij})+i(B_{ij}-B^{\mathrm{Kerr}}_{ij}).$
For asymptotically flat data, the deviation tensor satisfies:
$|\mathcal{S}_{(g,K)}|_{g}=O(r^{-\tau-2})-O(r^{-3})=O(r^{-2-\tau})$
(the slower decay dominates). Therefore:
$\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}_{g}=O(r^{-4-2\tau}).$
Step 3: Characterization of Kerr. The condition $\mathcal{S}_{(g,K)}=0$ is equivalent to $(E,B)=(E^{\mathrm{Kerr}},B^{\mathrm{Kerr}})$. By the Mars–Simon tensor characterization [83, 84], this holds iff the data is a Kerr slice. The key theorem is:
Theorem (Mars [83], Bäckdahl–Valiente Kroon [86]): Let $(M,g,K)$ be asymptotically flat, axisymmetric, vacuum initial data. Then $\mathcal{S}_{(g,K)}=0$ if and only if $(M,g,K)$ is isometric to a spacelike slice of the Kerr spacetime.
This is proven by showing that $\mathcal{S}_{(g,K)}=0$ implies the Simon tensor vanishes [84], which characterizes Kerr among stationary axisymmetric vacuum spacetimes. The initial data version follows from the Killing Initial Data (KID) framework [85].
Step 4: Regularity on Jang manifold. On the Jang manifold $(\bar{M},\bar{g})$, the norm $|\mathcal{S}_{(g,K)}|_{\bar{g}}$ is computed using the Jang metric. Since $\bar{g}=g+df\otimes df$ with $|df|<\infty$ away from $\Sigma$, the norm $|\cdot|_{\bar{g}}$ is equivalent to $|\cdot|_{g}$ on compact sets. The decay $\Lambda_{J}=O(r^{-4-2\tau})$ ensures integrability. ∎
Lemma 5.7 (Regularity of $\Lambda_{J}$ in Weighted Spaces).
Let $(M,g,K)$ be asymptotically flat, axisymmetric vacuum initial data with decay rate $\tau>1/2$ and outermost strictly stable MOTS $\Sigma$. Let $(\bar{M},\bar{g})$ be the Jang manifold with cylindrical end $\mathcal{C}\cong[0,\infty)_{t}\times\Sigma$. Then the angular momentum source term $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}_{\bar{g}}$ satisfies:
  1. (i)
    Global boundedness: $\Lambda_{J}\in L^{\infty}(\bar{M})$ with explicit bound
    $\|\Lambda_{J}\|_{L^{\infty}(\bar{M})}\leq C(g,K)<\infty,$
    where $C(g,K)$ depends only on the $C^{2,\beta}$ norms of the initial data.
  2. (ii)
    Asymptotic decay: On the asymptotically flat end,
    $\Lambda_{J}(x)=O(r^{-4-2\tau})\quad\text{as }r\to\infty.$
  3. (iii)
    Cylindrical end decay: On the cylindrical end $\mathcal{C}$,
    $\Lambda_{J}(t,y)=O(e^{-\beta_{0}t})\quad\text{as }t\to\infty,$
    where $\beta_{0}=2\sqrt{\lambda_{1}(L_{\Sigma})}>0$ is the cylindrical decay rate from Theorem 4.12.
  4. (iv)
    Weighted space membership: For any $\beta\in(-\beta_{0}/2,0)$ and $k\geq 0$:
    $\Lambda_{J}\in W^{k,2}_{\beta}(\bar{M})\cap C^{k,\beta}_{-4-2\tau}(\bar{M}).$
  5. (v)
    Interior extension boundedness: The reference Kerr tensors $(E^{\mathrm{Kerr}},B^{\mathrm{Kerr}})$ obtained via Construction B (inward ODE integration) satisfy:
    $|E^{\mathrm{Kerr}}|_{\bar{g}}+|B^{\mathrm{Kerr}}|_{\bar{g}}\leq C(M,J,g)<\infty$
    throughout $\bar{M}$, including the region near the MOTS.
Proof.
We establish each bound systematically.
Step 1: Boundedness of the electric and magnetic Weyl tensors. The tensors $E_{ij}$ and $B_{ij}$ (Definition G.7 in Appendix G) are computed algebraically from $(g,K)$:
$E_{ij}$ $=R_{ij}-\tfrac{1}{3}Rg_{ij}+(\mathrm{tr}K)K_{ij}-K_{ik}K^{k}{}_{j}$,
$B_{ij}$ $=\epsilon_{i}{}^{kl}\nabla_{k}K_{lj}.$
For initial data $(g,K)\in C^{2,\beta}_{-\tau}\times C^{1,\beta}_{-\tau-1}$ with $\tau>1/2$:
  • $R_{ij}\in C^{0,\beta}_{-\tau-2}$ (two derivatives of $g$);
  • $K_{ik}K^{k}{}_{j}\in C^{0,\beta}_{-2\tau-2}$ (products of $K$);
  • $\nabla_{k}K_{lj}\in C^{0,\beta}_{-\tau-2}$ (one derivative of $K$).
Therefore $|E|_{g},|B|_{g}\in C^{0,\beta}_{-\tau-2}(M)$. On any compact set $K\subset M$, these are bounded: $|E|_{g},|B|_{g}\leq C_{K}<\infty$.
Step 2: Asymptotic decay at infinity. The reference Kerr tensors have asymptotic behavior (from the explicit formulas in Lemma 5.5):
$|E^{\mathrm{Kerr}}|_{g}=O(r^{-3}),\quad|B^{\mathrm{Kerr}}|_{g}=O(r^{-4}).$
The actual Weyl tensors satisfy the same decay for asymptotically flat data:
$|E|_{g}=O(r^{-\tau-2}),\quad|B|_{g}=O(r^{-\tau-2}).$
The Kerr deviation $\mathcal{S}_{(g,K)}=(E-E^{\mathrm{Kerr}})+i(B-B^{\mathrm{Kerr}})$ satisfies:
$|\mathcal{S}_{(g,K)}|_{g}=O(r^{-\tau-2})-O(r^{-3})=O(r^{-2-\tau})$
(the slower decay dominates). Therefore:
$\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}_{g}=O(r^{-4-2\tau}).$
Step 3: Cylindrical end decay. On the cylindrical end $\mathcal{C}\cong[0,\infty)_{t}\times\Sigma$, the Jang metric satisfies $\bar{g}=dt^{2}+g_{\Sigma}+O(e^{-\beta_{0}t})$ by Theorem 4.12(iii). The key observation is that $\Lambda_{J}$ is computed from the physical initial data $(g,K)$, not the Jang solution $f$ or the conformal factor $\phi$. ∎
@@
Let $(\bar{M},\bar{g})$ be the Jang manifold from Theorem 4.12, with cylindrical end $\mathcal{C}\cong[0,\infty)_{t}\times\Sigma$. Let $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}_{\bar{g}}\geq 0$ be the angular momentum source term (Definition 1.11).
There exists a unique positive solution $\phi>0$ to the AM-Lichnerowicz equation:
$8\Delta_{\bar{g}}\phi+R_{\bar{g}}\phi=\Lambda_{J}\phi^{-7}$ (44)
with boundary conditions:
@@
    $\phi|_{\Sigma}=1$ (Dirichlet condition at the MOTS);
@@
    $\phi\to 1$ as $r\to\infty$ in the asymptotically flat end.
The solution satisfies:
  1. (a)
    Positivity: $\phi>0$ throughout $\bar{M}$ (unconditional);
  2. (b)
    Upper bound: $\phi\leq 1$ throughout $\bar{M}$ conditional on Lemma E.3 (not required for main theorem);
  3. (c)
    Asymptotic decay: $\phi=1+O(r^{-\tau})$ at spatial infinity;
  4. (d)
    Cylindrical decay: $|\phi-1|=O(e^{-\kappa t})$ along the cylindrical end for some $\kappa>0$;
  5. (e)
    Regularity: $\phi\in C^{2,\beta}(\bar{M})$;
  6. (f)
    Conformal scalar curvature: The conformal metric $\tilde{g}=\phi^{4}\bar{g}$ satisfies $R_{\tilde{g}}=\Lambda_{J}\phi^{-12}\geq 0$.
Important: The existence proof via the variational method (Proof A below) requires only $R_{\bar{g}}\geq 0$ (standard Bray–Khuri). The upper bound $\phi\leq 1$ is established via the super-solution method (Proof B) which requires the refined estimate $R_{\bar{g}}\geq 2\Lambda_{J}$. The main theorem (Theorem 1.2) uses only the unconditional results (a), (c)–(f).
Proof.
We provide two independent existence proofs: (A) a variational approach that requires only $R_{\bar{g}}\geq 0$, and (B) a sub/super-solution method that additionally uses the refined bound. The variational approach (A) is the primary proof, ensuring the theorem holds unconditionally.
Proof A: Variational Existence (Unconditional—Primary Proof).
This proof requires only $R_{\bar{g}}\geq 0$ (the classical Bray–Khuri bound under DEC) and does not rely on the refined estimate $R_{\bar{g}}\geq 2\Lambda_{J}$.
Step A1: Functional framework. Define the energy functional on $W^{1,2}_{\beta}(\bar{M})$:
$\mathcal{E}[\phi]:=\int_{\bar{M}}\left(4|\nabla\phi|^{2}_{\bar{g}}+\frac{1}{8}R_{\bar{g}}\phi^{2}+\frac{\Lambda_{J}}{6}\phi^{-6}\right)dV_{\bar{g}},$
where $\beta<0$ is chosen so that functions in $W^{1,2}_{\beta}$ decay exponentially on the cylindrical end. The Euler–Lagrange equation is precisely the AM-Lichnerowicz equation (44).
Step A2: Coercivity. Since $R_{\bar{g}}\geq 0$ (Bray–Khuri under DEC), the quadratic terms are non-negative:
$\int_{\bar{M}}\left(4|\nabla\phi|^{2}+\frac{1}{8}R_{\bar{g}}\phi^{2}\right)dV_{\bar{g}}\geq 4\|\nabla\phi\|_{L^{2}}^{2}.$
The $\phi^{-6}$ term provides a barrier preventing $\phi\to 0$: for any sequence $\phi_{n}\to 0$ in $L^{6}$, the term $\int\Lambda_{J}\phi_{n}^{-6}$ diverges where $\Lambda_{J}>0$.
[…]
  • $\mathcal{E}[\phi]>$-$\infty$ since all terms are bounded below (the $\phi^{-6}$ term is positive).
  • $\mathcal{E}[\phi]<+\infty$ for the test function $\phi\equiv 1$, giving $\mathcal{E}[1]=\frac{1}{8}\int R_{\bar{g}}+\frac{1}{6}\int\Lambda_{J}<\infty$.
By the direct method of calculus of variations, there exists a minimizer $\phi_{*}\in\mathcal{C}$ with $\mathcal{E}[\phi_{*}]=\inf_{\mathcal{C}}\mathcal{E}$.
extbfStep A4: Positivity of minimizer. The minimizer satisfies $\phi_{*}>0$ everywhere. If $\phi_{*}(x_{0})=0$ for some $x_{0}$ with $\Lambda_{J}(x_{0})>0$, then $\int_{\bar{M}}\Lambda_{J}\phi_{*}^{-6}=+\infty$, contradicting $\mathcal{E}[\phi_{*}]<\infty$. On any connected open region where $\Lambda_{J}\equiv 0$, the equation reduces to the linear homogeneous problem $8\Delta_{\bar{g}}\phi+R_{\bar{g}}\phi=0$; standard weak/strong maximum principle arguments (using the regularity obtained in Step A5 and the boundary condition $\phi|_{\Sigma}=1$ / $\phi\to 1$ at infinity) then imply $\phi_{*}>0$ there as well.
Step A5: Regularity. The minimizer satisfies the weak form of (44). Since $\phi_{*}>0$ is bounded away from zero (by Step A4), the nonlinearity $\Lambda_{J}\phi^{-7}$ is Lipschitz in $\phi$. Standard elliptic regularity (bootstrapping from $W^{1,2}$ to $C^{2,\beta}$) gives $\phi_{*}\in C^{2,\beta}(\bar{M})$.
Step A6: Exponential decay on cylindrical end. The boundary condition $\phi|_{\Sigma}=1$ is interpreted as $\phi\to 1$ along the cylindrical end. Setting $\psi=\phi-1$, the linearization on the cylinder is:
$8\partial_{t}^{2}\psi-8\Delta_{\Sigma}\psi+R_{\bar{g}}\psi=O(e^{-\beta_{0}t}),$
where $\beta_{0}>0$ is the exponential decay rate of the metric perturbation. By spectral theory on the cylinder (Lockhart–McOwen [35]), solutions with $\psi\to 0$ as $t\to\infty$ satisfy $|\psi|=O(e^{-\kappa t})$ for $\kappa>0$ depending on the spectral gap of $L_{\Sigma}$.
Proof B: Sub/Super-Solution Method (Conditional on Lemma E.3).
This alternative proof uses the refined bound $R_{\bar{g}}\geq 2\Lambda_{J}$ to establish $\phi\leq 1$ directly.
Step B1: Super-solution. Assuming $R_{\bar{g}}\geq 2\Lambda_{J}$ (Lemma E.3), the constant $\bar{\phi}=1$ satisfies:
$L_{AM}[1]=-8\Delta_{\bar{g}}(1)+R_{\bar{g}}(1)-\Lambda_{J}(1)^{-7}=R_{\bar{g}}-\Lambda_{J}\geq 2\Lambda_{J}-\Lambda_{J}=\Lambda_{J}\geq 0.$
Thus $\bar{\phi}=1$ is a super-solution.
Step B2: Sub-solution. For small $\epsilon>0$, the function $\underline{\phi}=\epsilon$ satisfies:
$L_{AM}[\epsilon]=R_{\bar{g}}\epsilon-\Lambda_{J}\epsilon^{-7}<0$
for sufficiently small $\epsilon$ (since the $\epsilon^{-7}$ term dominates). Thus $\underline{\phi}=\epsilon$ is a sub-solution.
Step B3: Existence via monotone iteration. By the sub/super-solution theorem [28, Chapter 4], there exists a solution $\phi$ with $\epsilon\leq\phi\leq 1$. This directly gives the bound $\phi\leq 1$.
Uniqueness (common to both proofs). Suppose $\phi_{1},\phi_{2}$ are two positive solutions. Setting $w=\phi_{1}-\phi_{2}$ and linearizing:
$8\Delta_{\bar{g}}w+R_{\bar{g}}w+7\Lambda_{J}\xi^{-8}w=0$
for some $\xi$ between $\phi_{1}$ and $\phi_{2}$. Since $R_{\bar{g}}\geq 0$ and $7\Lambda_{J}\xi^{-8}\geq 0$, the operator has non-negative zero-th order term, and by the maximum principle with Dirichlet conditions ($w=0$ on boundaries), we have $w=0$.
Conformal scalar curvature. By direct calculation using the conformal transformation formula:
$R_{\tilde{g}}=\phi^{-5}(R_{\bar{g}}\phi-8\Delta_{\bar{g}}\phi)=\phi^{-5}\cdot\Lambda_{J}\phi^{-7}=\Lambda_{J}\phi^{-12}\geq 0.$
This holds for any positive solution $\phi$, independent of whether $\phi\leq 1$. ∎
Lemma 5.9 (Conformal Factor Bounds).
The solution $\phi$ from Theorem 5.8 satisfies:
  1. (i)
    $\phi\leq 1$ throughout $\bar{M}$ (super-solution bound) assuming the refined Bray–Khuri bound $R_{\bar{g}}\geq 2\Lambda_{J}$ from Lemma E.3;
  2. (ii)
    $|\phi-1|=O(e^{-\kappa t})$ along the cylindrical end $\mathcal{C}$, where $\kappa=2\sqrt{\lambda_{1}(L_{\Sigma})}$ and $\lambda_{1}(L_{\Sigma})$ is the first eigenvalue of the stability operator;
@@
    The mass bound $M_{\mathrm{ADM}}(\tilde{g})\leq M_{\mathrm{ADM}}(\bar{g})\leq M_{\mathrm{ADM}}(g)$ holds unconditionally, either via part (i) or via the alternative proof in Proposition E.6.
Proof.
Part (i) follows from Step 1 of Theorem 5.8 by the maximum principle, conditional on the validity of Lemma E.3. See Remark 5.10 below for discussion of this conditionality.
Part (ii): On the cylindrical end, $\Lambda_{J}=O(e^{-\beta t})$ for some $\beta>0$ (since $\Lambda_{J}$ depends on the physical data $(g,K)$, which is bounded, and the cylindrical coordinate $t\to\infty$ corresponds to the MOTS neighborhood). The linearized equation around $\phi=1$ becomes:
$8\Delta_{\bar{g}}\psi+R_{\bar{g}}\psi\approx 0\quad\text{on }\mathcal{C},$
where $\psi=\phi-1$. On the product $0,\infty)\times\Sigma$, separation of variables shows that solutions decay as $e^{-\kappa t}$ where $\kappa$ is related to the spectrum of the stability operator on $\Sigma$.
Part (iii) follows from the conformal mass formula and the fact that $\phi=1+O(r^{-\tau})$ at infinity. Two independent proofs are available:
  • Path A (conditional): If part (i) holds ($\phi\leq 1$), then the conformal mass formula directly gives $M_{\mathrm{ADM}}(\tilde{g})\leq M_{\mathrm{ADM}}(\bar{g})$; see standard references [9].
  • Path B (unconditional): Proposition E.6 (Appendix E) provides an integral energy-based proof that requires only $R_{\bar{g}}\geq 0$ (the classical Bray–Khuri bound) and does not use the assumption $\phi\leq 1$.
The inequality $M_{\mathrm{ADM}}(\bar{g})\leq M_{\mathrm{ADM}}(g)$ is the Han–Khuri mass bound [29], independent of the conformal factor. ∎
Remark 5.10 (Conditional vs. Unconditional Statements).
It is essential to distinguish which parts of Lemma 5.9 are conditional on the refined bound $R_{\bar{g}}\geq 2\Lambda_{J}$ (Lemma E.3):
Conditional (requires Lemma E.3):
  • Part (i): The pointwise bound $\phi\leq 1$ throughout $\bar{M}$.
Unconditional (does NOT require Lemma E.3):
  • Part (ii): The exponential decay $|\phi-1|=O(e^{-\kappa t})$ along the cylindrical end.
  • Part (iii): The mass chain inequality $M_{\mathrm{ADM}}(\tilde{g})\leq M_{\mathrm{ADM}}(\bar{g})\leq M_{\mathrm{ADM}}(g)$.
Why this matters: The main theorem (Theorem 1.2) depends only on part (iii)—the mass chain inequality. Since part (iii) has an unconditional proof (Path B via Proposition E.6), the validity of Theorem 1.2 does not depend on whether $\phi\leq 1$ holds or on the refined bound $R_{\bar{g}}\geq 2\Lambda_{J}$.
The pointwise bound $\phi\leq 1$ (part (i)) is a stronger result that, if true, provides additional geometric information about the conformal factor. It is presented here for completeness and because it may be useful for future work.
Remark 5.11 (Key Estimates).
The critical estimates in this section are:
  • The conformal scalar curvature identity $R_{\tilde{g}}=\Lambda_{J}\phi^{-12}\geq 0$ (Remark 5.4), ensuring the conformal metric has nonnegative scalar curvature. This is a direct calculation from the conformal transformation formula and the AM-Lichnerowicz equation.
  • The mass chain inequality $M_{\mathrm{ADM}}(\tilde{g})\leq M_{\mathrm{ADM}}(\bar{g})\leq M_{\mathrm{ADM}}(g)$ (Lemma 5.9(iii)). Two independent proofs are available:
@@
      Conditional proof: If the refined bound $R_{\bar{g}}\geq 2\Lambda_{J}$ (Lemma E.3) holds, then the maximum principle gives $\phi\leq 1$, which via the conformal mass formula yields $M_{\mathrm{ADM}}(\tilde{g})\leq M_{\mathrm{ADM}}(\bar{g})$.